OpenStudy (anonymous):

SOLVE BY COMPLETING THE SQUARE : X^2+6X+3=19

jimthompson5910 (jim_thompson5910):

x^2+6x+3=19 x^2+6x=19-3 x^2+6x=16 x^2+6x+9=16+9 (x+3)^2=25 x+3 = +-sqrt(25) x+3 = sqrt(25) or x+3 = -sqrt(25) x+3 = 5 or x+3 = -5 x = 5-3 or x = -5 - 3 x = 2 or x = -8

OpenStudy (anonymous):

X^2+6X+3=19 ===> X^2+6X+9 =19+3+9 ===> (x+3)^2=31

OpenStudy (anonymous):

WHOSE RIGHT JIM OR RAHEEN

jimthompson5910 (jim_thompson5910):

my answer is correct and has been verified by wolfram alpha

jimthompson5910 (jim_thompson5910):

http://www.wolframalpha.com/input/?i=x^2%2B6x%2B3%3D19

OpenStudy (anonymous):

OKAYY

OpenStudy (anonymous):

SO THE ANSWER IS -8

OpenStudy (anonymous):

(x + 3)^2 + 3 - 9 = 19 (x+3)^2 = 19 -3+9 = 25 (x+ 3) = 5 or -5 x = 2 or -8

OpenStudy (anonymous):

two answers 2 and -8

OpenStudy (ybarrap):

jim is right, 2 & -8, two answers